OpenStudy (anonymous):

HOW can i solve this? f(x) = x^2 + 3 g(x) = x- 7 (g o f) (2) PLEASE SHOW THE SOLUTION. THANKS! >:D<

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

http://www.twiddla.com/563194

OpenStudy (anonymous):

What's the matter? Afraid of a little help?

OpenStudy (saifoo.khan):

Hero!

OpenStudy (anonymous):

pls help

OpenStudy (anonymous):

f(x) = x^2 + 3 g(x) = x- 7 (g o f) (2) =g(f(2)) f(2)=2^2+3=7 g(f(2))=g(7)=7-7 =0

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

ywc